Clarke has a controlling interest in Rogers's outstanding stock. At the current year-end, the following information has been accumulated for these two companies:

  

Separate Operating Income

Dividends Paid

  Clarke

$552,500                   

$95,000      

(includes a $96,000 net unrealized gross profit on intra-entity inventory transfers)

  Rogers

335,000                   

55,000      


Assume that Clarke owns 80 percent of Rogers's voting stock, but the companies elect to file separate tax returns. What is the total amount of income taxes that these two companies pay for the current period?
Answer #1

The parent owns 80% of the subsidiary's voting stock and the companies have elected to file separate tax returns. In this scenario, the intra-entry dividends are not taxable since both companies qualify as an affiliated group. The income tax payable for both companies is found from multiplying the tax rate by their operating income

Tax (clarke ) = 552500 * 40% = 221000

Tax (Rogers ) = $335000* 40% = 134000

* assuming tax rate = 40%